Use Runge-Kutta method of order 2 to solve y′ = xy, y(1) = 1, in [1, 1.4] by taking step-length h = 0.2


1
Expert's answer
2021-02-02T05:16:58-0500

Runge - Kutta method of order 2

We begin with two functions evaluates of the form:

"F_1 = hf(x,y)"

"F_2 = hf(x+\alpha h, y + \beta F_1)"

The "\alpha" and "\beta" are unknown quantities. The idea was to take a linear combination of the

"F_1 \space and \space F_2" terms to obtain an approximation for the y value at "x = x_o+h," and to find appropriate values of "\alpha" and "\beta".

By comparing the values obtains using Taylor's Series method and the above terms (I will spare you the details here), they obtained the following, which is Runge-Kutta Method of Order 2:

"y(x+h) = y(x) + \frac{1}{2}(F_1+F_2)"

We have: "y'(x, y) = xy" y(1) = 1 [1, 1.4] h = 0.2

We start with x = 0 and y = 1. We'll find the F values first.

"F_1 = hf(x,y) = 0.2*1*1=0.2"

"x + \frac{h}{2} = 1 + 0.1 = 1.1"

"y + \frac{F_1}{2} = 1+0.1 = 1.1"

"y(x+h) = y(x) + \frac{1}{2}(F_1+F_2) = 1*1+\frac{1}{2}(1.1+1.1) = 2.1"